Solve the following LPP by the two-phase simplex method.

 Max Z = x1 + x2 − x3

 Subject to

4x1 + x2 + x3 = 4

3x1 + 2x2 - x4 = 6

x1,x2,x3 ≥ 0


1
Expert's answer
2021-06-14T16:21:50-0400

Solution:

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate

1. As the constraint-1 is of type '=' we should add artificial variable A1

2. As the constraint-2 is of type '=' we should add artificial variable A2

3. As the constraint-3 is of type '≤' we should add slack variable S1

4. As the constraint-4 is of type '≤' we should add slack variable S2

5. As the constraint-5 is of type '≥' we should subtract surplus variable Sand add artificial variable A3

After introducing slack,surplus,artificial variables:

Max Z=-A1-A2-A3 subject to

"4x_1+x_2+x_3+A_1=4
\\
3x_1+2x_2+A_2=6
\\
x_1+S_1=0
\\
x_2+S_2=0
\\x_3-S_3+A_3=0"

and 

x1,x2,x3,S1,S2,S3,A1,A2,A3≥0



Since all Zj-Cj≥0

Hence, optimal solution has arrived with the value of variables as :

x1=0,x2=0,x3=4

Max Z=0

But this solution is not feasible

because the final solution violates the 2nd constraint  3 x1 + 2 x2    = 6.

and the artificial variable A2 appears in the basis with a positive value of 6.

So phase-2 is not possible.